OpenStudy (anonymous):

Help Solving: |x+2|=3 I don't know what the lines mean, step by step please.

jimthompson5910 (jim_thompson5910):

|x| = k means x = k or x = -k for some positive number k

jimthompson5910 (jim_thompson5910):

so |x+2|=3 means x+2 = 3 or x+2 = -3

jimthompson5910 (jim_thompson5910):

solve each of the last two equations for x